OpenStudy (anonymous):

How many gallons of a 3% salt solution must be mixed with 50 gallons of a 7% solution to obtain a 5% solution?

OpenStudy (anonymous):

x*3+50*7=(50+x)*5 3x+350=250+5x 2x=100 x=50 50 gallons

OpenStudy (anonymous):

Sorry I'm just having real trouble with these problems but isn't the 3% turned into a .03 decimal. I checked the back of my book and that is right btw.

OpenStudy (anonymous):

yes... .03*x+.07*50=.05(x+50) 3*x+7*50=5(x+50) multiplying with 100

OpenStudy (anonymous):

Thank you.